OpenStudy (anonymous):

How many years older will you be 1.60 billion seconds from now? (Assume a 365-day year.)

OpenStudy (mathstudent55):

\( 1.60 \times 10^9 ~sec \times \dfrac{1~min}{60~sec} \times \dfrac{1~hr}{60~min} \times \dfrac{1~day}{24~hour} \times \dfrac{1~year}{365~day} \)

OpenStudy (anonymous):

is that 10 to the 9th power? 10^9

OpenStudy (mathstudent55):

Yes. 10^9 is one billion.

OpenStudy (anonymous):

thank you!

OpenStudy (mathstudent55):

You're welcome. Did you get an answer in years?

OpenStudy (anonymous):

yes